under e stellata ieast frequently found in wine fermentations. Its sugar metabolism is being studied atasts in continuous culture with fructose (CsHi20s) as the carbon source and ammonium Estequiometria microbiana-Coeficiente Respirométrico fen source, the yield of biomass from fructose is 0.025 g g, the yield of ethanol nom uctose is 0.21 g g, and the yield of glycerol (CsH&Os) from fructose is 0.07 gg aerobic conditions C HeO) from i a) If the rate of co b) What is the RQ for this culture? gar constion is 190 g h what is the rate of oxygen consumption in g h 7

Solution, please.
Answer #1

a) from the problem gram per gram 0.025g biomass, 0.21g ethanol, 0.07g glycerol rest 0.695g co2 should be formed

for 180g of fructose 192g of oxygen is required, for 190 g of fructose theoritically 202g of o2 is required

theoritically from 180g of fructose 264g of co2 will be liberated, here some of the carbon is converting in to biomass and other products. for 190g 202g of o2 is required. but yeast liberating per gram of fructose 0.695g of co2 hence 0.695*264 = 183.48g co2

b) RQ is c02/o2 = 182.48/264 = 0.95